OpenStudy (anonymous):

Can someone Write and solve a linear system to find the cost of renting a double-occupancy room and the cost of renting a single-occupancy room. (See Q in comments)

OpenStudy (anonymous):

A hotel rents a double-occupancy room for $30 more than a single-occupancy room. One night, the hotel took in $3115 after renting 15 double-occupancy rooms and 26 single-occupancy rooms.

OpenStudy (anonymous):

single: $65 double: $90

OpenStudy (anonymous):

$95 my bad

OpenStudy (anonymous):

Can you write a linear equation for it too please ?

OpenStudy (anonymous):

I'm not really good at that but...here's the work let single room as 'S' and double room as 'S+30'. So, 15(S+30) + 26(S) = 3115 41(S) + 15(30) = 3115 41(S) + 450 = 3115 41(S) = 3115 - 450 = 2665 S = 2665/41 = 65. So single room cost = 65. And double room cost = 65 + 30 = 90.

OpenStudy (anonymous):

thanks . :)